A normal distribution has the mean of 150.5. if 70.54% of the area under the curve lies to the left of 177.5 find a. the standard deviation?​





1
Expert's answer
2021-11-09T13:22:33-0500

"P(x<177.5)=0.7054"


"z=0.54"


"z=\frac{x-\mu}{\sigma}"


the standard deviation:


"\sigma=\frac{x-\mu}{z}=\frac{177.5-150.5}{0.54}=50"